# WakenLLM: Evaluating Reasoning Potential and Stability in LLMs via Fine-Grained Benchmarking

### v4

@@ V U V U V U V U
Table 7: The percentage of LLMs demonstrating the Vague Perception phenomenon among six LLMs. Split by two different root causes: (1) V: Verifiable samples with T/F labels, but the model output Unknown. (2) U: Unverifiable Samples with Unknown label, but the model output Unknown simply because of incapacity rather than fully understand this is objectively unsolvable.
Search-Based Planning Frameworks
Instead of generating a single chain, Tree-of-Thought (ToT) views partial solutions as nodes in a tree and performs breadth-first search with value heuristics (Yao et al., 2023b). Decomposed Prompting turns a complex task into callable sub-prompts that can be solved (and verified) independently, achieving strong zero-shot gains on compositional reasoning benchmarks (Khot et al., 2023). Reasoning-via-Planning (RAP) couples an LLM world-model with Monte-Carlo Tree Search for long-horizon tasks (Hao et al., 2023). Cochain (Chain-of-Collaboration Prompting) extends this line by orchestrating multiple lightweight agents whose specialized sub-chains are merged via a knowledge-graph plan, outperforming single-agent CoT on workflow-style tasks while using fewer tokens (Zhao et al., 2025).
Appendix C Related Work Details
Chain-of-Thought (CoT) prompting (Wei et al., 2022a) first showed that furnishing step-wise exemplars markedly improves LLM performance on arithmetic and commonsense problems. Follow-up work refines this paradigm: scratchpads (Nye et al., 2021), least-to-most prompting (Zhou et al., 2022), and self-consistency (Wang et al., 2023c). More recently, multi-agent and symbolic reasoning approaches like multiagent debate (Du et al., 2024), symbolic chain-of-thought (Xu et al., 2024), and neurosymbolic reasoning with logic provers (Olausson et al., 2023) further strengthen factuality and logical consistency.
Tool-Augmented Reasoning
Program-Aided LMs (Gao et al., 2022) delegate arithmetic or logical computation to Python. ReAct (Yao et al., 2022) interleaves reasoning and external tools. CRITIC (Gou et al., 2024) validates intermediate answers with execution tools. Chain-of-Verification (Dhuliawala et al., 2024) reduces hallucinations by verifying sub-claims against evidence.
Iterative Self-Refinement and Critique
Self-Refine (Madaan et al., 2023) establishes a generate–criticise–rewrite loop. Ambiguity handling and uncertainty decomposition are also explored: AmbigQA (Min et al., 2020), Decomposing Uncertainty (Hou et al., 2024), and rank-based calibration (Huang et al., 2024).
Datasets for Reasoning and Verification
Classic benchmarks like BoolQ (Clark et al., 2019), SQuAD 2.0 (Rajpurkar et al., 2018), FEVER (Thorne et al., 2018), and HotpotQA (Yang et al., 2018) stress-test fact verification, ambiguity, and multi-hop reasoning.
Active Prompting
Active-Prompt (Diao et al., 2024) selects informative exemplars to boost performance under budget constraints. Meanwhile, Ling et al. (2025) denotes different prompt coverage leads to various effects.
Faithfulness and Safety
Beyond correctness, methods ensure interpretability and reliability, such as Faithful-CoT (Lyu et al., 2023), symbolic verification (Xu et al., 2024), and multi-agent debate (Du et al., 2024).
Appendix D Formula Details
D.1 False and Unknown Converting
As demonstrated in 2.1, we denote the notion of TCR. Likewise, we define the concepts regarding false converting and unknown converting similarly:
$FC_{f}^{n}\;=\;\sum_{y_{f}\in D_{\mathrm{VP}}^{n}}\hskip-6.99997pt\mathbf{1}\bigl(\hat{y}_{f}^{n}\neq y_{f},\;\hat{y}_{f}\notin U\bigr)$
$FCR_{f}^{n}\;=\;\frac{\bigl|FC_{f}^{n}\bigr|}{\bigl|D_{\mathrm{VP}}^{n}\bigr|}$
$UC_{f}^{n}\;=\;\sum_{y_{f}\in D_{\mathrm{VP}}^{n}}\hskip-6.99997pt\mathbf{1}\bigl(\hat{y}_{f}^{n}\in U,\;y_{f}\notin U\bigr)$
$UCR_{f}^{n}\;=\;\frac{\bigl|UC_{f}^{n}\bigr|}{\bigl|D_{\mathrm{VP}}^{n}\bigr|}$
Note that, Unknown Converting (UC) is still categorized into Vague Perception phenomenon: Outputting unknown when facing verifiable samples. We can tell the following principles for any stage, n belongs to 1 or 2:
$UCR_{f}^{n}+FCR_{f}^{n}+TCR_{f}^{n}\;=\;100\%$
$UC_{f}^{n}\cup FC_{f}^{n}\cup TC_{f}^{n}\;=\;D_{VP}^{n}$
$UC_{f}^{n}\cap FC_{f}^{n}\cap TC_{f}^{n}\;=\;\emptyset$
Our metrics systematically define and categorize the flowing of output distribution by means of above metrics.
D.2 Overall Converting & Latent Ability
We clarify the relationship between Overall Converting Rate (OCR) in Section 2.1 with the latent ability in 4.5, as we depicted, by comparing True Converting samples (TC) with $D_{VP}$, we can get a empirical ability of how many samples can be eventually converted to right labels within all Vague Perception samples, whereas the latter indicates overall accuracy gain among the all samples D in a broader aspect. By choosing different objects to compare with, we can gain insights not only in the phenomenon alone, but also the general LLMs’ reasoning potential.
@@ D.2 Overall Converting & Latent Ability
Table 8: Detailed Remind-then-Guide (RtG) setting of GPT-3.5 and GPT-4o. M denotes Misguiding Rate: the percentage of input labels are replaced with wrong one. We break down $TCR^{1}$ into verifiable and unverifiable data: $TCR_{v}^{1}$ and $TCR_{u}^{1}$
@@ V U V U
50 14.65 10.70 5.99 45.00
0 17.75 15.21 11.27 60.00
Table 9: Detailed Remind-then-Guide (RtG) setting of LLaMA-3.1-8B and LLaMA-3.1-405B. M denotes Misguiding Rate: the percentage of input labels are replaced with wrong one. We break down $TCR^{1}$ into verifiable and unverifiable data: $TCR_{v}^{1}$ and $TCR_{u}^{1}$. "/" denotes there is no input for stage 2, as false converted samples ($FC^{1}$ from stage 1 don’t exist.
@@
50 4.40 6.71 0.26 15.00
0 5.09 10.42 0.00 56.00
Table 10: Detailed Remind-then-Guide (RtG) setting of Qwen-2.5-7B and Qwen-2.5-72B. M denotes Mis-guiding Rate: the percentage of input labels are replaced with wrong one. We break down $TCR^{1}$ into verifiable and unverifiable data: $TCR_{v}^{1}$ and $TCR_{u}^{1}$
@@
50 21.09 4.30 13.10 28.00
0 24.22 9.38 17.42 45.00
Table 11: Six LLMs’ $TCR^{1}$ and $TCR^{2}$ break down into verifiable and unverifiable samples: $TCR_{v}^{1}$, $TCR_{u}^{1}$, $TCR_{v}^{2}$, $TCR_{u}^{2}$, among all four datasets.
@@
GPT-3.5 FLD 12.61 0 9.28 5.15
FOLIO 14.41 0 4.21 37.89
Science 21.05 0 3.33 0
Arts 22.92 3.12 4.00 0
@@
Science 44.98 0.96 5.95 0
Arts 18.81 5.45 8.57 0
LLaMA-3.1-405B FLD 5.43 15.22 0 53.73
FOLIO 9.59 19.17 0 29.92
Science 10.23 22.62 2.50 0
Arts 4.17 0.23 0 0
Qwen2.5-7B FLD 47.19 1.12 0 40.00
@@
Science 15.83 0 10.00 0
Arts 23.31 5.93 25.37 0
Qwen2.5-72B FLD 46.39 1.81 0 9.80
@@
Science 10.81 0 3.12 0
Arts 18.75 6.25 9.76 0

### v5

@@ V U V U V U V U
Table 8: The percentage of LLMs demonstrating the Vague Perception phenomenon among six LLMs. Split by two different root causes: (1) V: Solvable samples with T/F labels, but the model output Unknown. (2) U: Unsolvable Samples with Unknown label, but the model output Unknown simply because of incapacity rather than fully understand this is objectively unsolvable.
Appendix I Detailed TCR Analysis
Our TCR (True Converting Rate) analysis via different stages can offer valuable insights on model selections in multi-turn frameworks, as multi-agent and multi-dialogue workflows are more and more widely used now. Table 12 decomposes $TCR^{1}$ and $TCR^{2}$ into stages regarding (V) and unsolvable (U) samples. GPT-3.5 exhibits moderate first stage converting on solvable samples (e.g., $12.61\%$ on FLD and $14.41\%$ on FOLIO) but few converting rate on unsolvable examples ($TCR^{1}_{u}$). Its second stage, however, can salvage many unverifiable cases on FOLIO, yielding a $TCR^{2}_{u}$ of $37.89\%$. GPT-4o delivers stronger $TCR^{1}$ on solvable samples in story-based datasets like FOLIO ($33.73\%$) and shows significant second stage gains on unsolvable samples ($44.03\%$ on FOLIO and $26.87\%$ on FLD). This suggests that GPT-4o is more suited for workflows requiring extended dialogues to resolve ambiguous inputs.
LLaMA-3.1-8B stands out with consistently high $TCR^{1}_{s}$ across all four datasets, reaching $44.20\%$ on FOLIO and $44.98\%$ on Science. In contrast, the 405B model counterpart attains lower $TCR^{1}_{s}$ and relies on stage 2 converting to correct the mistakes; for instance, on FLD it has $TCR^{1}_{s}$ of only $5.43\%$ but $TCR^{2}_{u}$ of $53.73\%$. This implies that smaller LLaMA models can resolve most solvable queries in a single turn, whereas the 405B model might need deeper chains of reasoning.
Among Qwen series, the 7B model achieves the highest $TCR^{1}_{s}$ on FLD ($47.19\%$) and Science ($15.83\%$), but its stage 2 converting occur mostly on unsolvable samples; for example, $TCR^{2}_{u}$ reaches $40.00\%$ on FLD and $35.09\%$ on FOLIO. The 72B model exhibits a more balanced behavior, with moderate $TCR^{1}_{s}$ and lower $TCR^{2}_{u}$, indicating that it resolves solvable samples earlier and requires fewer subsequent turns. Taken together, these findings underscore the importance of considering both stages: models with high $TCR^{1}_{s}$ are advantageous for straightforward, solvable samples, whereas those with substantial $TCR^{2}_{u}$ may better handle ambiguous or mislabeled inputs in multi-turn dialogues.
Appendix J Vague Perception Details
J.1 Evaluations
Detailed distribution of the Vague Perception samples is shown in Table 8. We can tell even though it is prevalent across different datasets and models, there is a general inclination. Among all extracted samples, models will larger parameters are more likely to output Unknown, falling into the second part. With a smaller parameter it is very likely to be the opposite. For Qwen series, the distribution can be relatively stable across two models, however, LLaMA-3.1 and GPT series both demonstrate stronger differences within the same model family.
J.2 Causes
Even though MoE structure models like GPT-4o demonstrates superior generalization ability across different fields, however, the generalization ability across different presentation styles is still unstable.
Appendix K Terminology Clarification
Abstention behavior is widely researched, such reasoning biases combined with knowledge boundary is rather important to offer trustworthy and unbiased outputs. We further elaborate the introduced the term Vague Perception to distinguish the differences between existing concepts.
Abstention
Abstention typically indicates LLMs themselves know there are malicious, poison elements within the prompt and refuse to answer.
Knowledge Boundary
Knowledge Boundary means LLMs know when to answer, and when not to, to prevent hallucination and promote abstention when being asked something beyond the ability.
Hallucinations
Hallucination is about including non-exist facts or reasoning in the output, the paradigm of how this is induced is still controversial, but it is mainly because of detailed prompt asking and multi-turn dialogue.
Vague Perception (Ours)
Given an unknown option, LLMs are able to answer correctly. However, they are potentially inclined to output unknown, even if the problem can be solved within the their reasoning ability.
Appendix L Important Findings Wrap-up
L.1 Unknown Label, Stronger Conformity
Previous conformity research mainly focuses on multi-turn dialogue workflows zhu2025conformitylargelanguagemodels. However, containing uncertainty elements is a fundamental part of conformity. We mainly consider this is because of imbalanced training data source that leads to relatively vulnerable resistance against misleading.
L.2 Correct Prediction, Flawed Reasoning
As experiments shown, CGR and RPC results are mainly negative, we can conclude that most LLMs predicted correct labels with flawed reasoning processes. We attribute this to the essence of reinforcement learning (RL) algorithms being a result-oriented methodology, the algorithm only cares about the final prediction, rather than intermediate processes, casting potential risks in trustworthiness.
L.3 Bigger Parameter Sizes, More Cautious
Our comparisons show that for both GPT, LLaMA and Qwen series, LLMs with higher parameters are inclined to output unknown. Similar to Degradation analysis in Section 5.7, this is a unified conclusion obtained in multiple experiments. Opposite to the findings in (weng2025dothinkconformitylarge), where models are stable and more resistant against misguiding, as we include the Unknown option within the prompt, models with bigger parameter sizes like to conform and stick to this label, we attribute this to Over-Cautious. This is exactly opposite to the findings in (DBLP:journals/nature/ZhouSMMFH24), casting unstable and random biases exhibiting under various scenarios.
L.4 LLMs Selection in Multi-stage Workflows
As addressed in Section 5.1 and Appendix I, LLaMA series demonstrate superior performances at early stages, it is recommended to be placed at front parts. However, for GPT-4o and Qwen series, the sequence of placing is highly dependent on the dataset forms, they exhibit distinct yet related behaviors between Story-based and Fact-based datasets. For example, GPT-4o demonstrates better stage 1 performance in Fact-based datasets, so we suggest putting it prior to other models when dealing with samples with atomic statements, and the opposite dealing with samples containing long passages; Whereas Qwen series show relatively balanced performance in Story-based datasets, and superior ability at stage 1 in Fact-based datasets.
Our experiments and strategies are purely based on sequential workflows, and parallel workflows are not tested or used in our work.
Appendix M Pseudo-Code Pipelines Visualization
As shown in Algorithm 12 and 3, the pipeline of different experimental settings within our pipeline is demonstrated below. Differences from Vanilla settings are marked orange.
0: Dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ with labels $y_{i}\in\{\mathrm{T},\mathrm{F},\mathrm{U}\}$;base LLM model $\mathcal{M}$; two stimulation prompts $P_{1}$ and $P_{2}$
0: Stage-1 categories $\mathrm{TC}^{1},\mathrm{UC}^{1},\mathrm{FC}^{1}$ and Stage-2 categories $\mathrm{TC}^{2},\mathrm{FC}^{2}$; reasoning store $\mathcal{R}$
[…]
2:for all $(x_{i},y_{i})$ in $\mathcal{D}$ do
3:  $\hat{y}_{i}\leftarrow\mathcal{M}(x_{i})$ // first-pass prediction
4:  if ( $y_{i}\in\{\mathrm{T},\mathrm{F}\}$ and $\hat{y}_{i}=\mathrm{U}$ ) or ( $y_{i}=\mathrm{U}$ and $\hat{y}_{i}=\mathrm{U}$ and model lacks reasoning ) then
5:   $\mathrm{D_{VP}}\leftarrow\mathrm{D_{VP}}\cup\{(x_{i},y_{i})\}$ // collect Vague Perception samples
6:  end if
7:end for
[…]
10:for all $(x_{i},y_{i})$ in $\mathrm{D_{VP}}$ do
11:  $(\hat{y}_{i}^{(1)},r_{i})\leftarrow\mathcal{M}(\text{stimulation}(x_{i},P_{1}))$ // first-stage stimulation returns both predicted label and reasoning process $r_{i}$
12:  $\mathcal{R}[i]\leftarrow r_{i}$ // store reasoning process for sample $i$
13:  if $\hat{y}_{i}^{(1)}=y_{i}$ then
14:   $\mathrm{TC}^{1}\leftarrow\mathrm{TC}^{1}\cup\{(x_{i},y_{i})\}$ // True Converting at stage 1
15:  else if $\hat{y}_{i}^{(1)}=\mathrm{U}$ then
16:   $\mathrm{UC}^{1}\leftarrow\mathrm{UC}^{1}\cup\{(x_{i},y_{i})\}$ // Unknown Converting at stage 1
17:  else
18:   $\mathrm{FC}^{1}\leftarrow\mathrm{FC}^{1}\cup\{(x_{i},y_{i})\}$ // False Converting at stage 1
19:  end if
20:end for
[…]
22:for all $(x_{i},y_{i})$ in $\mathrm{FC}^{1}$ do
23:  $\hat{y}_{i}^{(2)}\leftarrow\mathcal{M}(\text{stimulation}(x_{i},P_{2},\mathcal{R}[i]))$ // second-stage stimulation: input the $FC^{1}$ along with its previous reasoning $r_{i}$
24:  if $\hat{y}_{i}^{(2)}=y_{i}$ then
25:   $\mathrm{TC}^{2}\leftarrow\mathrm{TC}^{2}\cup\{(x_{i},y_{i})\}$ // Corrected at stage 2
26:  else
27:   $\mathrm{FC}^{2}\leftarrow\mathrm{FC}^{2}\cup\{(x_{i},y_{i})\}$ // Still incorrect at stage 2
28:  end if
29:end for
30:return $\mathrm{TC}^{1},\mathrm{UC}^{1},\mathrm{FC}^{1},\mathrm{TC}^{2},\mathrm{FC}^{2},\mathcal{R}$
Algorithm 2 Pipeline of the Remind-then-Guide (RtG) Setting.
0: Vague Perception set $\mathrm{D_{VP}}$ and Stage-1 outputs $\{(\hat{y}_{i}^{(1)},r_{i}^{(1)})\}_{i\in\mathrm{D_{VP}}}$ obtained by Algorithm 1
0: Ablated Stage-1 converting statistics over $T$ rounds
1: // Pre-processing, Vague Perception extraction and the first Stage-1 stimulation
2: // follow Algorithm 1 and are omitted here.
[…]
4:for $t=2$ to $T$ do
5:  $\mathrm{TC}^{1,(t)},\mathrm{UC}^{1,(t)},\mathrm{FC}^{1,(t)}\leftarrow\emptyset$ // Stage-1 categories in round $t$
6:  for all $(x_{i},y_{i})$ in $\mathrm{D_{VP}}$ do
7:   $(\hat{y}_{i}^{(t)},r_{i}^{(t)})\leftarrow\mathcal{M}(\text{stimulation}(x_{i},P_{1},\mathcal{R}[i]))$ // re-use Stage-1 prompt but remind with previous reasoning $\mathcal{R}[i]$
8:   if $\hat{y}_{i}^{(t)}=y_{i}$ then
9:    $\mathrm{TC}^{1,(t)}\leftarrow\mathrm{TC}^{1,(t)}\cup\{(x_{i},y_{i})\}$
10:   else if $\hat{y}_{i}^{(t)}=\mathrm{U}$ then
11:    $\mathrm{UC}^{1,(t)}\leftarrow\mathrm{UC}^{1,(t)}\cup\{(x_{i},y_{i})\}$
12:   else
13:    $\mathrm{FC}^{1,(t)}\leftarrow\mathrm{FC}^{1,(t)}\cup\{(x_{i},y_{i})\}$
14:   end if
15:   $\mathcal{R}[i]\leftarrow r_{i}^{(t)}$ // update stored reasoning to the latest round
16:  end for
[…]
18:end for
19:return $\{\mathrm{TC}^{1,(t)},\mathrm{UC}^{1,(t)},\mathrm{FC}^{1,(t)}\}_{t=1}^{T}$, $\mathcal{R}$
Algorithm 3 Ablation Study of the Remind-then-Guide (RtG) Setting.
@@ Appendix I Detailed TCR Analysis
Table 9: Detailed Remind-then-Guide (RtG) setting of GPT-3.5 and GPT-4o. We break down $TCR^{1}$ into solvable and unsolvable samples: $TCR_{s}^{1}$ and $TCR_{u}^{1}$. M: Misguiding Rate
@@
50.00 14.65 10.70 5.99 45.00
0.00 17.75 15.21 11.27 60.00
Table 10: Detailed Remind-then-Guide (RtG) setting of LLaMA-3.1-8B and LLaMA-3.1-405B. We break down $TCR^{1}$ into solvable and unsolvable samples: $TCR_{s}^{1}$ and $TCR_{u}^{1}$. "/" denotes there is no input for stage 2, as false converted samples ($FC^{1}$) from stage 1 don’t exist. M: Misguiding Rate.
@@
50.00 4.40 6.71 0.26 15.00
0.00 5.09 10.42 0.00 56.00
Table 11: Detailed Remind-then-Guide (RtG) setting of Qwen-2.5-7B and Qwen-2.5-72B. We break down $TCR^{1}$ into solvable and unsolvable data: $TCR_{s}^{1}$ and $TCR_{u}^{1}$. M: Misguiding Rate.
@@
50.00 21.09 4.30 13.10 28.00
0.00 24.22 9.38 17.42 45.00
Table 12: Six LLMs’ $TCR^{1}$ and $TCR^{2}$ break down into solvable and unsolvable samples: $TCR_{s}^{1}$, $TCR_{u}^{1}$, $TCR_{s}^{2}$, $TCR_{u}^{2}$, among all four datasets.
@@ V U V U
GPT-3.5 FLD 12.61 0.00 9.28 5.15
FOLIO 14.41 0.00 4.21 37.89
Science 21.05 0.00 3.33 0.00
Arts 22.92 3.12 4.00 0.00
@@
Science 44.98 0.96 5.95 0.00
Arts 18.81 5.45 8.57 0.00
LLaMA-3.1-405B FLD 5.43 15.22 0.00 53.73
FOLIO 9.59 19.17 0.00 29.92
Science 10.23 22.62 2.50 0.00
Arts 4.17 0.23 0.00 0.00
Qwen2.5-7B FLD 47.19 1.12 0.00 40.00
@@
Science 15.83 0.00 10.00 0.00
Arts 23.31 5.93 25.37 0.00
Qwen2.5-72B FLD 46.39 1.81 0.00 9.80
@@
Science 10.81 0.00 3.12 0.00
Arts 18.75 6.25 9.76 0.00
Appendix N Prompt Templates & Examples
@@
Baseline 1 Template: Direct Reasoning Prompt Template: Here is the Hypothesis: {conclusion} Now These are the ’facts’: {facts} Please carefully evaluate the relationship between the facts and the hypothesis. Return UNKNOWN if the facts are insufficient to make a definitive conclusion. Return __PROVED__ only if the facts support the hypothesis. Return __DISPROVED__ only if the facts contradict the hypothesis. Output Conclusion: __PROVED__, __DISPROVED__, or __UNKNOWN__. Input & Output Example: Facts: All birds have feathers. Penguins are birds. Feathers provide insulation. Conclusion: Penguins can maintain their body temperature. LLM Reasoning Output: Penguins are birds, all birds have feathers, and feathers provide insulation. Therefore, penguins can maintain their body temperature. Label Output: __PROVED__ True Answer: __PROVED__
Baseline 2 Template: Chain of Thought (CoT) Prompt Template: Here is the Hypothesis: {conclusion} Now These are the ’facts’: {facts} Please carefully evaluate the relationship between the facts and the hypothesis. Let’s think step by step. Return UNKNOWN if the facts are insufficient to make a definitive conclusion. Return __PROVED__ only if the facts support the hypothesis. Return __DISPROVED__ only if the facts contradict the hypothesis. Output Conclusion: __PROVED__, __DISPROVED__, or __UNKNOWN__. Input & Output Example: Facts: If it rains, the ground gets wet. If the ground gets wet, plants grow better. It is raining today. Conclusion: Plants will grow better today. LLM Reasoning Output: Step 1: It is raining today (given fact). Step 2: If it rains, the ground gets wet (given fact). Step 3: Therefore, the ground gets wet today (from Step 1 and Step 2). Step 4: If the ground gets wet, plants grow better (given fact). Step 5: Therefore, plants will grow better today (from Step 3 and Step 4). Label Output: __PROVED__ True Answer: __PROVED__
Baseline 3 Template: Few-shot Learning Prompt Template: Example 1: Hypothesis: Water boils at 100°C. Facts: At standard atmospheric pressure, water boils at 100°C. Output: __PROVED__ Example 2: Hypothesis: The sun rises in the west. Facts: The sun rises in the east every day. Output: __DISPROVED__ Here is the Hypothesis: {conclusion} Now These are the ’facts’: {facts} Please carefully evaluate the relationship between the facts and the hypothesis. Return UNKNOWN if the facts are insufficient to make a definitive conclusion. Return __PROVED__ only if the facts support the hypothesis. Return __DISPROVED__ only if the facts contradict the hypothesis. Output Conclusion: __PROVED__, __DISPROVED__, or __UNKNOWN__. Input & Output Example: Facts: All mammals are warm-blooded. Dogs are mammals. Warm-blooded animals maintain constant body temperature. Conclusion: Dogs maintain constant body temperature. LLM Reasoning Output: Following the examples, dogs are mammals, all mammals are warm-blooded, and warm-blooded animals maintain constant body temperature. Therefore, dogs maintain constant body temperature. Label Output: __PROVED__ True Answer: __PROVED__
Baseline 4 Template: Tree of Thought (ToT) Prompt Template (Multi-step Process): Step 1: Generate Multiple Reasoning Paths Here is the Hypothesis: {conclusion} Now These are the ’facts’: {facts} Generate 3 different reasoning paths to evaluate this case. Each path should consider different aspects or perspectives: Step 2: Evaluate Each Path Original Question: {original_prompt} Reasoning Path: {thought} Rate this reasoning path from 1-10 (1=very poor, 10=excellent) and explain why: Step 3: Select Best Path and Generate Final Answer Based on this reasoning path: {best_thought} Original Question: {original_prompt} Provide the final answer: __PROVED__, __DISPROVED__, or __UNKNOWN__ Input & Output Example: Facts: All metals conduct electricity. Copper is a metal. If something conducts electricity, it can be used in wires. Conclusion: Copper can be used in wires. LLM Reasoning Output: Path 1: Copper is a metal → All metals conduct electricity → Copper conducts electricity → Can be used in wires. (Score: 9/10) Path 2: Consider material properties → Copper has conductivity → Suitable for wires. (Score: 7/10) Path 3: Direct application → Copper wires exist → Therefore can be used. (Score: 6/10) Best Path: Path 1. Final Answer: __PROVED__ Label Output: __PROVED__ True Answer: __PROVED__
Baseline 5 Template: Plan & Solve Prompt Template (Two-step Process): Step 1: Create Plan Here is the Hypothesis: {conclusion} Now These are the ’facts’: {facts} Before solving this problem, let’s create a detailed step-by-step plan: 1. First, we need to… 2. Then, we should… 3. Finally, we will… Please create a comprehensive plan for evaluating this case: Step 2: Execute Plan Original Question: {original_prompt} Plan: {plan} Now let’s execute this plan step by step. Follow the plan carefully and provide the final answer: __PROVED__, __DISPROVED__, or __UNKNOWN__ Input & Output Example: Facts: All plants need sunlight to grow. Trees are plants. If something needs sunlight, it requires light energy. Conclusion: Trees require light energy. LLM Reasoning Output: Plan: 1. Identify that trees are plants. 2. Apply the rule that all plants need sunlight. 3. Connect sunlight to light energy. 4. Conclude that trees require light energy. Execution: Step 1: Trees are plants (given). Step 2: All plants need sunlight (given). Step 3: Therefore, trees need sunlight. Step 4: Sunlight provides light energy, so trees require light energy. Label Output: __PROVED__ True Answer: __PROVED__
Baseline 6: Self-Reflection (Step 1 - Initial Answer Generation) Step 1: Initial Answer Generation Prompt Template For Dataset1/2 (Facts-based Datasets): Here is the Hypothesis: {conclusion} Now These are the ’facts’: {facts} Please carefully evaluate the relationship between the facts and the hypothesis. Return UNKNOWN if the facts are insufficient to make a definitive conclusion. Return __PROVED__ only if the facts support the hypothesis. Return __DISPROVED__ only if the facts contradict the hypothesis. Output Conclusion: __PROVED__, __DISPROVED__, or __UNKNOWN__. Please provide your initial answer with detailed step-by-step reasoning. Think carefully about the relationship between the facts and the hypothesis/answer. For Dataset3/4 (Story-based Datasets): Question: {question} Facts: {facts} Answer: {correct_answer} Based on the facts provided, determine if they can: 1. Prove the answer is correct (return __PROVED__) 2. Disprove the answer (return __DISPROVED__) 3. Are insufficient to determine the answer’s correctness (return __UNKNOWN__) Output Format: Only the Label: __PROVED__, __DISPROVED__, or __UNKNOWN__. Please provide your initial answer with detailed step-by-step reasoning. Think carefully about the relationship between the facts and the hypothesis/answer.
Baseline 6: Self-Reflection (Step 2 - Reflection) Step 2: Reflection Prompt Template Original Question: {original_prompt} Your Initial Answer: {initial_answer} Now, carefully reflect on your answer: 1. Are there any logical errors or gaps in your reasoning? 2. Did you consider all the given facts? 3. Is there any evidence that contradicts your conclusion? 4. Are there alternative interpretations you missed? 5. Is your confidence level justified? Provide a critical reflection on your initial answer, pointing out any potential issues or improvements:
Baseline 6: Self-Reflection (Step 3 - Refinement) Step 3: Refinement Prompt Template Original Question: {original_prompt} Your Initial Answer: {initial_answer} Your Reflection: {reflection} Based on your reflection, provide a refined final answer. If your initial answer was correct, confirm it. If there were issues, provide the corrected answer. Final Answer: __PROVED__, __DISPROVED__, or __UNKNOWN__